OpenStudy (anonymous):

As of today, Sam has read one-fifth of a 540-page book and Michael has read one-fourth of the same book. Starting tomorrow, Sam will read another one-tenth of the book each day and Michael will read another one-fifteenth each day. After how many days will Sam have read more pages than Michael?

OpenStudy (anonymous):

\[\text{Today Sam read}\frac{1}{5}(540) = \text{108 pages,and Michael read}\\ \frac{1}{5}(540) = 135 \text{pages.}\\\\540-( 108+135)=297\ \text{pages left}\\\text{After that, the proportion is :} \frac{3}{2} \ \text{, so the answer is 3 days} \]